# Learning-Order Autoregressive Models with Application to Molecular Graph Generation

### v1

@@ Abstract
Autoregressive models (ARMs) have become the workhorse for sequence generation tasks, since many problems can be modeled as next-token prediction. While there appears to be a natural ordering for text (i.e., left-to-right), for many data types, such as graphs, the canonical ordering is less obvious. To address this problem, we introduce a variant of ARM that generates high-dimensional data using a probabilistic ordering that is sequentially inferred from data. This model incorporates a trainable probability distribution, referred to as an order-policy, that dynamically decides the autoregressive order in a state-dependent manner. To train the model, we introduce a variational lower bound on the exact log-likelihood, which we optimize with stochastic gradient estimation. We demonstrate experimentally that our method can learn meaningful autoregressive orderings in image and graph generation. On the challenging domain of molecular graph generation, we achieve state-of-the-art results on the QM9 and ZINC250k benchmarks, evaluated using the Fréchet ChemNet Distance (FCD).
@@ 1 Introduction
  • We apply LO-ARMs to two molecule generation tasks (i.e., QM9 and ZINC250k) and obtain state-of-the-art results for both datasets, as measured by sample quality metrics such as Fréchet ChemNet Distance (FCD) (Preuer et al., 2018). We find that LO-ARMs can learn consistent orderings to generate new molecules with high validity and uniqueness.
@@ 2.2 Autoregressive Generation as Unmasking Process
Given that each discrete data dimension (or token) $x_{i},i\in\{1\dots L\}$ takes $m$ categorical values we further augment the space with an extra auxiliary category or mask. Thus, we represent each $x_{i}$ as an $m+1$-dimensional (rather than $m$-dimensional) one-hot vector where the final $m+1$-th value indicates that $x_{i}$ is masked. Then we can model the generation process with ARMs as an unmasking process. Specifically, starting with a fully masked state $\bar{\bm{x}}=(\bar{x}_{1},\ldots,\bar{x}_{L})$, where $\bar{x}_{i}$ represents the mask, at each step, we choose a dimension $\bar{x}_{i}$ and “unmask it” which means to sample a categorical value $x_{i}$ among the $m$ categories. Then we repeat this process until all dimensions are unmasked, which yields a final generated data point $\bm{x}$. Under this representation our model also connects with recent masked discrete diffusion models (see, e.g., Shi et al., 2024) as we further discuss in Related Work. However, note that unlike discrete diffusion models we do not specify a forward process in our framework, but only the unmasking or backward process.
@@ 3 Learning Order ARM
We extend AO-ARMs by modifying the prior uniform distribution $p(\sigma)$ to obtain a learnable distribution over orderings. We call this distribution the order-policy since it follows an autoregressive structure that dynamically decides the next dimension to generate by conditioning on the already generated data dimensions of $\bm{x}$. Next, we focus on the general description of our method and the key design choices, and delay the technical details of specifying it to practical tasks, including images (Appendix A) and molecular graphs (Appendix E) generation. Specifically, this section unfolds as follows. We introduce the order-policy in Section 3.1, define the learning objective in Section 3.2, and describe options for the model parametrizations and neural network architectures in Section 3.3.
@@ 3.2 Training with Variational Inference
$\frac{L}{2}\{\left(\nabla_{\theta}\log q_{\theta}(\bm{z}_{<i}^{2}|\bm{x})-\nabla_{\theta}\log q_{\theta}(\bm{z}_{<i}^{2}|\bm{x})\right)\Delta F$
@@
Once training is completed, we can generate new samples from the model using Algorithm 2. Specifically, we can generate a new sample $\hat{\bm{x}}=(\hat{x}_{1},\ldots,\hat{x}_{L})$ autoregressively starting from a fully masked state $\bar{\bm{x}}=(\bar{x}_{1},\ldots,\bar{x}_{L})$. To do this, at step $i\in\{1,\ldots,L\}$, we unmask $\bar{x}_{z_{i}}$ to $\hat{x}_{z_{i}}$, which is sampled with the categorical distribution. Note that we only need the model order-policy over $z_{i}\sim p_{\theta}(\cdot|\bm{z}_{<i},\bm{x}_{\bm{z}_{<i}})$ and the categorical distribution $\hat{x}_{z_{i}\sim}p_{\theta}(x_{z_{i}}|\bm{x}_{\bm{z}_{<i}})$ during inference.
@@ 3.3 Parametrization of the Distributions
Finally, we have already described the general form of the variational factors $q_{\theta}(z_{i}=k|\bm{z}_{<i},\bm{x}_{\bm{z}_{<i}})$ in Equation 12, so what remains is to specify the vector of logits $g_{\theta}(\bm{x})$. We explore two parametrizations: 1) A shared-torso parameterization where we add another final linear layer to $f_{\theta}$ (similarly to (ii) above) outputting $L$ values for $g_{\theta}(\bm{x})$. 2) A separate NN with $L$ outputs to model $g_{\theta}(\bm{x})$. The motivation of this latter option is to more freely capture the form of the variational distribution.
@@ 5 Evaluation and Analysis
We evaluate LO-ARM on two widely-adopted molecular graph generation benchmasks: QM9 and ZINC250K (Xu et al., 2019; Dwivedi et al., 2023). We follow the standard setup - e.g., in (Eichelsbacher & Reinert, 2008; Vignac et al., 2023; Jo et al., 2022), including data preprocessing, network parametrization, and evaluation metrics. We represent molecules as graphs with dense adjacency matrices (see Appendix E). For each model variant, we generate $16384$ samples and evaluate them according to two aspects: 1) validity and uniqueness of individual molecules and 2) Frèchet ChemNet Distance (FCD), which evaluates the distance between the distributions of true and generated molecules using the activations of ChemNet (Preuer et al., 2018). Note that, these two classes of metrics are not explicitly correlated, since invalid molecules may have latent activations close to valid ones. As we can see in Table 2, the validity of methods with similar FCDs may vary in a big range (e.g., GraphAF, GraphARM, EDP-GNN and SPECTRE). On the other hand, if the true dataset has high quality, we could still expect molecule validity to be improved when just optimizing FCD-related objectives, e.g., the lower bound of log-likelihoods. In this study, we focus more on the evaluation with FCD to show the generative capability of LO-ARM, and we will show how our validity and uniqueness correlate with FCD.
@@
We observe that there is generally a gap between the validity of ARMs and that of the best method on this task. We hypothesize that this is caused by the errors accumulated during autoregressive sampling, as we unmask a new dimension at a time without refining the previously unmasked. Still, as demonstrated by the gain in validity compared to AO-ARM and GraphARM, LO-ARM has largely fixed this problem by training an order-policy that maximizes the ELBO. Moreover, this accumulated error can be further reduced by employing Top-p sampling (Holtzman et al., 2019), a popular method used in autoregressive language models to address similar problems. Specifically, the Top-$p$ sampler only draws from dimensions with high probability, and $p$ corresponds to the cumulative probability of the top dimensions ranked with respect to their probabilities. Specifically, $p=1.0$ regresses to sampling the entire distribution, and the lower the $p$, the greedier the sampler. We tested this approach by sharpening the distribution of the classifier $f_{\theta}$ after the policy has sampled a dimension. After incorporating the Top-$p$ sampler, we can see that both the validity of Biased-AO-ARM-topp (top-$p=0.9$) and LO-ARM-st-st-topp (top-$p=0.9$) increase by a large margin, while at a cost of downgraded FCDs. As expected, the Top-$p$ sampling introduces a small bias, shifting the model’s distribution away from the true data distribution. We provide further ablation studies on this effect in Appendix D.
@@
Finally, we visualize the generation ordering learned by LO-ARM on ZINC250k in Figure 1 and Section B.2. Without imposing any inductive bias on orderings, LO-ARM learns to favour an edge-first ordering, i.e, 1) building skeleton of a molecule with chemical bonds, 2) infilling the rest of the adjacency matrix no_edge states or the “imaginery” bonds, and finally 3) infilling atoms (See Appendix B.2). This is different from the node-first ordering proposed in Kong et al. (2023). To analyze the consistency of this learned ordering, we sample $30,000$ molecules with three LO-ARM-st-sep models trained with different random seeds, and find that $99.9\%$ of the molecules are generated with the edge-first ordering (See Appendix C). Note that learning a context-dependent ordering (that varies across different data/molecules) within the edge or node dimensions is still critical to performance, as we can see by comparing LO-ARM with Biased-AO-ARM which just uniformly samples within edge or node dimensions.
@@ 6 Conclusions and Discussion
We have introduced LO-ARM, a novel autoregressive method that can learn the ordering for generating data without canonical ordering. We derive a simple variational lower bound which can be optimized with unbiased stochastic gradients. In addition, we design a scalable training algorithm which makes LO-ARM easily adaptable to real-world tasks. Moreover, we evaluate LO-ARM on two popular molecular generation tasks, and achieve state-of-the-art results in terms of FCD, and competitive results on molecule validity and uniqueness. In particular, we show that LO-ARM can learn a consistent ordering for generating new molecules without requiring inductive bias on generation orderings, which in turn contributes to a significant performance gain for generative modeling.
@@ E.1 Data Representation
We follow the standard setup (Vignac et al., 2023; Eijkelboom et al., 2024) to transform raw data represented as SMILES strings to dense graphs, $\mathbf{G}=(\mathbf{H}_{n},\mathbf{H}_{e},\mathbf{H}_{m})$, where $\mathbf{H}_{n}$ is the node vector, $\mathbf{H}_{e}$ is the edge adjacency matrix and $\mathbf{H}_{m}$ is the node mask. In particular, the node masks pad molecules with variable sizes to a static size, such that they can be processed with mini-batches. Moreover, same as DiGress, to represent the edge connections with a dense adjacency matrix, an “imaginery” bond is introduced, which we also call no_edge. This imaginary bond has no chemical effect when building molecules but helps us distinguish from the entries masked different masks. Now we introduce the masks used in our implementation.

### v2

@@ Abstract
Autoregressive models (ARMs) have become the workhorse for sequence generation tasks, since many problems can be modeled as next-token prediction. While there appears to be a natural ordering for text (i.e., left-to-right), for many data types, such as graphs, the canonical ordering is less obvious. To address this problem, we introduce a variant of ARM that generates high-dimensional data using a probabilistic ordering that is sequentially inferred from data. This model incorporates a trainable probability distribution, referred to as an order-policy, that dynamically decides the autoregressive order in a state-dependent manner. To train the model, we introduce a variational lower bound on the log-likelihood, which we optimize with stochastic gradient estimation. We demonstrate experimentally that our method can learn meaningful autoregressive orderings in image and graph generation. On the challenging domain of molecular graph generation, we achieve state-of-the-art results on the QM9 and ZINC250k benchmarks, evaluated across key metrics for distribution similarity and drug-likeless.
@@ 1 Introduction
  • We apply LO-ARMs to two molecule generation tasks (i.e., QM9 and ZINC250k) and obtain state-of-the-art results for both datasets, as measured by sample quality metrics such as Frèchet ChemNet Distance (FCD) (Preuer et al., 2018), Synthetic Accessibility Score (SAS), Quantitative Estimate of Drug-likeness (QED), and sample diversity.
  • We find that LO-ARMs can learn consistent orderings to generate new molecules with high validity and uniqueness.
@@ 2.2 Autoregressive Generation as Unmasking Process
Given that each discrete data dimension (or token) $x_{i},i\in\{1,\dots,L\}$ takes $m$ categorical values we further augment the space with an extra auxiliary category or mask. Thus, we represent each $x_{i}$ as an $m+1$-dimensional (rather than $m$-dimensional) one-hot vector where the final $m+1$-th value indicates that $x_{i}$ is masked. Then we can model the generation process with ARMs as an unmasking process. Specifically, starting with a fully masked state $\bar{\bm{x}}=(\bar{x}_{1},\ldots,\bar{x}_{L})$, where $\bar{x}_{i}$ represents the mask, at each step, we choose a dimension $\bar{x}_{i}$ and “unmask it” which means to sample a categorical value $x_{i}$ among the $m$ categories. Then we repeat this process until all dimensions are unmasked, which yields a final generated data point $\bm{x}$. Under this representation our model also connects with recent masked discrete diffusion models (see, e.g., Shi et al., 2024) as we further discuss in Related Work. However, note that unlike discrete diffusion models we do not specify a forward process in our framework, but only the unmasking or backward process.
@@ 3 Learning Order ARM
We replace the uniform prior distribution $p(\sigma)$ in AO-ARMs by with a learnable distribution over orderings. We call this distribution the order-policy since it dynamically decides the next dimension to generate by conditioning on the already generated data dimensions of $\bm{x}$. Next, we focus on the general description of our method and the key design choices, and delay the technical details of specifying it to practical tasks, including images (Appendix A) and molecular graphs (Appendix E) generation. Specifically, this section unfolds as follows. We introduce the order-policy in Section 3.1, define the learning objective in Section 3.2, and then describe options for the model parametrizations and neural network architectures in Section 3.3.
@@ 3.2 Training with Variational Inference
$\frac{L}{2}\{\left(\nabla_{\theta}\log q_{\theta}(\bm{z}_{<i}^{1}|\bm{x})-\nabla_{\theta}\log q_{\theta}(\bm{z}_{<i}^{2}|\bm{x})\right)\Delta F$
@@
Once training is completed, we can generate new samples from the model using Algorithm 2. Specifically, we can generate a new sample $\hat{\bm{x}}=(\hat{x}_{1},\ldots,\hat{x}_{L})$ autoregressively starting from a fully masked state $\bar{\bm{x}}=(\bar{x}_{1},\ldots,\bar{x}_{L})$. To do this, at step $i\in\{1,\ldots,L\}$, we unmask $\bar{x}_{z_{i}}$ to $\hat{x}_{z_{i}}$, which is sampled with the categorical distribution. Note that we only need the model order-policy over $z_{i}\sim p_{\theta}(\cdot|\bm{z}_{<i},\bm{x}_{\bm{z}_{<i}})$ and the categorical distribution $\hat{x}_{z_{i}}\sim p_{\theta}(\cdot|\bm{x}_{\bm{z}_{<i}})$ for inference.
@@ 3.3 Parametrization of the Distributions
Finally, we have already described the general form of the variational factors $q_{\theta}(z_{i}=k|\bm{z}_{<i},\bm{x})$ in Equation 12, so what remains is to specify the vector of logits $g_{\theta}(\bm{x})$. We explore two parametrizations: 1) A shared-torso parameterization where we add another final linear layer to $f_{\theta}$ (similarly to (ii) above) outputting $L$ values for $g_{\theta}(\bm{x})$. 2) A separate NN with $L$ outputs to model $g_{\theta}(\bm{x})$. The motivation of this latter option is to more freely capture the form of the variational distribution.
@@ 5 Evaluation and Analysis
We evaluate LO-ARM on two widely-adopted molecular graph generation benchmasks: QM9 and ZINC250K (Xu et al., 2019; Dwivedi et al., 2023). We follow the standard setup - e.g., in Eichelsbacher & Reinert (2008); Vignac et al. (2023); Jo et al. (2022), including data preprocessing, network parametrization, and evaluation metrics. We represent molecules as graphs with dense adjacency matrices (see Appendix E). For each model variant, we generate $16384$ samples and evaluate them according to two aspects: 1) validity and uniqueness of individual molecules and 2) Frèchet ChemNet Distance (FCD), which evaluates the distance between the distributions of true and generated molecules using the activations of ChemNet (Preuer et al., 2018). Note that, these two classes of metrics are not explicitly correlated, since invalid molecules may have latent activations close to valid ones. As we can see in Table 2, the validity of methods with similar FCDs may vary in a big range (e.g., GraphAF, GraphARM, EDP-GNN and SPECTRE). On the other hand, if the true dataset has high quality, we could still expect molecule validity to be improved when just optimizing FCD-related objectives, e.g., the lower bound of log-likelihoods. In this study, we focus more on the evaluation with FCD to show the generative capability of LO-ARM, and we will show how our validity and uniqueness correlate with FCD.
@@
Table 3: LO-ARM’s perforamance on chemistry-specific metrics on ZINC250k. Specifically, lower similarity indicates higher diversity for molecules in the dataset. On all these metrics, LO-ARM exceeds or matches the performance of the state-of-the-art results of the models. The results of other methods are cited from (Mazuz et al., 2023). V. is short for Validity.
[…]
Ground truth (test data) 0.75 2.76 0.35 100.0
JT-VAE 0.64 4.69 - 100.00
GCPN 0.65 4.53 - 99.00
Taiga 0.75 2.89 - 88.00
LO-ARM-st-sep (ours) 0.75 3.08 0.34 96.26
@@
We observe that there is generally a gap between the validity of ARMs and that of the best method on this task. We hypothesize that this is caused by the errors accumulated during autoregressive sampling, as we unmask a new dimension at a time without refining the previously unmasked. Still, as demonstrated by the gain in validity compared to AO-ARM and GraphARM, LO-ARM has largely fixed this problem by training an order-policy that maximizes the ELBO. Moreover, this accumulated error can be further reduced by employing Top-$p$ sampling (Holtzman et al., ), a popular method used in autoregressive language models to address similar problems. Specifically, the Top-$p$ sampler only draws from dimensions with high probability, and $p$ corresponds to the cumulative probability of the top dimensions ranked with respect to their probabilities. Specifically, $p=1.0$ regresses to sampling the entire distribution, and the lower the $p$, the greedier the sampler. We tested this approach by sharpening the distribution of the classifier $f_{\theta}$ after the policy has sampled a dimension. After incorporating the Top-$p$ sampler, we can see that both the validity of Biased-AO-ARM-topp (Top-$p=0.9$) and LO-ARM-st-st-topp (Top-$p=0.9$) increase by a large margin, while at a cost of downgraded FCDs. As expected, the Top-$p$ sampling introduces a small bias, shifting the model’s distribution away from the true data distribution. We provide further ablation studies on this effect in Appendix D.
@@
Thirdly, on the front of practical usefulness, the ultimate goal of the generative model is is to sample from the same chemical space that the training data comes from. On the other hand, the validity and uniqueness metrics are fairly saturated by existing methods, and the FCD on its own is not a strong enough predictor of molecule quality. To address this issue, we further evaluate LO-ARM-st-sep on the chemistry-specific metrics, including 1) Synthetic Accessibility Score (SAS), which measures the ease of synthesizing a chemical compound (i.e., the lower the score, the easier the compound to be synthesized), 2) Quantitative Estimate of Drug-likeness (QED) which evaluates how well a compound’s physicochemical properties align with those of successful administered drugs (the higher the score, the more aligned), and 3) pairwise similarity measures the diversity in samples (lower score corresponds to higher diversity). We visualize the distributions on these three metrics in Figure 2. On both datasets the distributions of these metrics calculated on the LO-ARM samples closely match the corresponding ground-truth (GT) data, a finding consistent with their favorable FCD scores. We also compare LO-ARM with other methods in the literature on these metrics. As shown in Table 3, LO-ARM exceeds or matches the previous best results in all three metrics.
Figure 2: Evaluating LO-ARM on (a) ZINC250k and (b) QM9) on the chemistry-specific metrics , i.e, Synthetic Accessibility Score (SAS), Quantitative Estimate of Drug-likeness (QED), and pairwise similarity between samples. On both datasets, the distributions of the samples generated with LO-ARM on these metrics are well-matched with the corresponding ground-truth (GT) data, which is well-aligned with their performance on FCD scores.
Finally, we visualize the generation ordering learned by LO-ARM on ZINC250k in Figure 1 and Section B.2. Without imposing any inductive bias on orderings, LO-ARM learns to favour an edge-first ordering, i.e, 1) building skeleton of a molecule with chemical bonds, 2) infilling the rest of the adjacency matrix “no-edge” states or the “imaginery” bonds, and finally 3) infilling atoms (See Appendix B.2). This is different from the node-first ordering proposed in Kong et al. (2023). To analyze the consistency of this learned ordering, we sample $30,000$ molecules with three LO-ARM-st-sep models trained with different random seeds, and find that $99.9\%$ of the molecules are generated with the edge-first ordering (See Appendix C). Note that learning a context-dependent ordering (that varies across different data/molecules) within the edge or node dimensions is still critical to performance, as we can see by comparing LO-ARM with Biased-AO-ARM which just uniformly samples within edge or node dimensions.
@@ 6 Conclusions and Discussion
We have introduced LO-ARM, a novel autoregressive method that can learn the ordering for generating data without canonical ordering. We derive a simple variational lower bound which can be optimized with unbiased stochastic gradients. In addition, we design a scalable training algorithm which makes LO-ARM easily adaptable to real-world tasks. Moreover, we evaluate LO-ARM on two popular molecular generation tasks, and achieve state-of-the-art results in terms of FCD and QED, and competitive results on molecule validity, uniqueness and SAS. In particular, we show that LO-ARM can learn a consistent ordering for generating new molecules without requiring inductive bias on generation orderings, which in turn contributes to a significant performance gain for generative modeling.
@@
Impact Statement
This paper presents the work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of this work, none of which we feel must be specifically highlighted here.
@@ E.1 Data Representation
We follow the standard setup (Vignac et al., 2023; Eijkelboom et al., 2024) to transform raw data represented as SMILES (Simplified Molecular Input Line Entry System) strings to dense graphs, $\mathbf{G}=(\mathbf{H}_{n},\mathbf{H}_{e},\mathbf{H}_{m})$, where $\mathbf{H}_{n}$ is the node vector, $\mathbf{H}_{e}$ is the edge adjacency matrix and $\mathbf{H}_{m}$ is the node mask. In particular, the node masks pad molecules with variable sizes to a static size, such that they can be processed with mini-batches. Moreover, same as DiGress, to represent the edge connections with a dense adjacency matrix, an “imaginery” bond is introduced, which we also call no_edge. This imaginary bond has no chemical effect when building molecules but helps us distinguish from the entries masked different masks. Now we introduce the masks used in our implementation.